OpenStudy (anonymous):

Simplify. Show your work. Someone please break it down and help, I am having troubles with this one! 5 1/3 +(-3 9/18)

OpenStudy (anonymous):

1 5/6

OpenStudy (anonymous):

rite @ybarrap

OpenStudy (anonymous):

ik i am but 4 back-up

OpenStudy (anonymous):

@ybarrap did you comment on this, if so it's not showing up :(

OpenStudy (anonymous):

i thot he was gonna @Cassie__Winfrey :(

OpenStudy (anonymous):

I neeeeeed helppppp on how to do thisssssss :(

OpenStudy (anonymous):

oh on how?

OpenStudy (anonymous):

The whole problem, I need it broken down so I can understand it

OpenStudy (anonymous):

first 9/18 = 1/2 ok ?

OpenStudy (ybarrap):

$$ 5 \cfrac{1}{3}+(-3 \cfrac{9}{18})\\ =\cfrac{16}{3}-\cfrac{63}{18}\\ =6\times \cfrac{16}{18}-\cfrac{63}{18}\\ =\cfrac{96}{18}-\cfrac{63}{18}\\ =\cfrac{96-63}{18}\\ =\cfrac{33}{18}\\ =\cfrac{3\times 11}{3\times 6}\\ =\cfrac{11}{6}\\ $$

OpenStudy (anonymous):

dafuq!^

OpenStudy (ybarrap):

Attaching, just in case you can't see that -- looks like Latex is broken

OpenStudy (anonymous):

Thank you so much! @ybarrap

OpenStudy (anonymous):

11/ 6 = 1 5/6

OpenStudy (ybarrap):

NP

OpenStudy (anonymous):

:(